[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Aug 25, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Aug 25 10:04:23 UTC 2017


HSU Shift Log for Fri Aug 25 2017 06:04:17 GMT-0400 (EDT)

Shift: A

Forecaster: L Avila

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: None

AWIPS-2/NCP issues: None

Other Techincal issues: No SHIPS for AL92 at 06Z (though EC SHIPS ran  
normally) Users occasionally getting "Access Denied" messages on NHC  
website. Reported problem to SDM. Appears to be a WOC problem, not an NHC  
problem. Waiting a minute and refreshing the page typically solves the  
problem.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170825/8a0c7715/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list